# Self-Improvement for Neural Combinatorial Optimization: Sample without Replacement, but Improvement

### v1

@@ 1 Introduction
Several learning-based methods have been proposed for solving combinatorial optimization (CO) problems in recent years. These neural approaches aim to enable a deep neural network to learn how to generate heuristics without manual crafting. A prominent paradigm in this direction is the constructive approach, where a solution to a problem instance is built step by step as a sequential decision problem. A neural network models the policy that guides these incremental decisions. It is trained by supervised learning (SL) using expert solutions as labeled data (Vinyals et al., 2015; Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022; Hottung et al., 2020; Drakulic et al., 2023; Luo et al., 2023) or by reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019b; Nazari et al., 2018; Chen & Tian, 2019; Ahn et al., 2020; d O Costa et al., 2020; Hottung & Tierney, 2020; Kwon et al., 2020; Ma et al., 2021; Kim et al., 2021; Park et al., 2021; Wu et al., 2021; Park et al., 2022; Zhang et al., 2020; Hottung et al., 2022; Zhang et al., 2024), which exploits the natural modeling of the sequential problem as a finite Markov decision process. While obtaining expert solutions from (exact) solvers can be expensive for SL-based methods, RL-based methods, especially policy gradient methods, can be challenging to train due to high memory requirements and strong hyperparameter sensitivity (Schulman et al., 2017; Henderson et al., 2018). Recent work attributes the poor ability of current methods to generalize to larger instances than those seen during training to the typically lightweight decoder of the used architectures (Drakulic et al., 2023; Luo et al., 2023). Increasing the size of the decoder structure can address this issue, but makes it impractical to train the model with policy gradient methods due to the increased computational cost.
In this paper, we move away from RL for neural CO towards a simple, problem-independent training scheme derived from the Cross-Entropy Method (CEM) (De Boer et al., 2005): We sample a set of solutions for randomly generated problem instances in each epoch using the current model. The sampled solution with the best objective function evaluation for each instance is treated as a pseudo-expert trajectory on which we can train the model in a supervised manner. By repeatedly ’self-imitating’ the best solutions found, we create an improving loop.
@@
The effectiveness of this strategy relies on the quality of the sampled solutions and the efficiency of the sampling process. In particular, we want to sample as few sequences as possible to get better and better solutions in each epoch. Simple repeated sampling from the conditional distribution given by the model at each sequential step can result in a set of sequences that lacks diversity, contains numerous duplicates, and requires a large number of samples to find a solution that enables the network to improve. This paper utilizes a sampling mechanism based on batch-wise Stochastic Beam Search (SBS) (Kool et al., 2019c) as introduced by (Shi et al., 2020). The mechanism draws samples without replacement in multiple rounds while maintaining a search tree. We suggest taking advantage of the batch- and round-wise mechanism to enhance its effectiveness. Given the sequences sampled in a single round, we estimate the expected value of the objective function and update the trie with the advantage of the sampled sequences, a strategy derived from a provable policy improvement operation. Furthermore, to balance the explore-exploit tradeoff, we couple this update strategy with Top-$p$ (nucleus) sampling (Holtzman et al., 2020). We gradually increase $p\leq 1$ after each round to allow more unreliable sequences. Our method applies to any constructive neural CO problem, unlike sampling strategies that leverage problem specifics to diversify and improve the sampled solutions (Kwon et al., 2020).
@@
(i) For constructive neural combinatorial optimization, we propose to sample diverse sequences without replacement from the current model and use the best sequences as pseudo-expert trajectories to imitate.
(ii) We draw the sequences in multiple rounds using batch-wise SBS. We develop a technique to make the underlying search tree more informed in each round by updating the sequence probabilities according to their advantage. Informally, we increase (decrease) the probability of better (worse) trajectories than expected. This update adds next to no computational overhead to the sampling method, and we derive this technique from a provable policy improvement.
(iii) We couple SBS with Top-$p$ sampling, where we start with $p<1$ and increase to $p=1$ over the rounds of sampling. While putting pressure on the model to disregard unreliable choices with a low $p$ works well in confident models, exploiting the model in such a way can hurt the performance in the early stages of training when the model itself is unreliable. By covering a range of $p$, we compromise between both cases.
(iv) We show on the TSP and CVRP that we can achieve comparable results to direct training on (near) optimal expert trajectories. We further present a pure self-attention architecture for JSSP that outperforms the current state of the art when trained with our method.
@@ 2 Related Work
Self-imitation in neural CO
The CEM (De Boer et al., 2005; Rubinstein, 1999) is a derivative-free optimization technique that can be summarized as generating a set of solution candidates, evaluating them according to some objective function, and selecting the best-performing candidates to guide the generation of new solutions. While this principle of pushing the policy toward better performing sampled solutions lies at the heart of the self-critical policy gradient methods, there is little work on directly cloning the behavior of trajectories sampled from the current model in neural CO (Bogyrbayeva et al., 2022; Bengio et al., 2021). In fact, independent of our work, Corsini et al. (2024) make the same observation. They present a ’self-labeling’ strategy that samples multiple solutions from the current policy and uses the best one as a pseudo-label for supervised training. However, they focus only on the JSSP and use vanilla sampling with replacement, in contrast to our main contribution, which updates the policy during the sampling without replacement process. We refer to Appendix E for a deeper comparison. Furthermore, Luo et al. (2023) show a hybrid problem-specific way to learn without any labeled data by pre-training a randomly initialized model with RL on small instances, followed by reconstructing partial solutions to refine an unlabeled training dataset. We show that we can outperform this strategy without relying on RL.
@@ 3 Learning Algorithm
In the following section, we formally set up the problem and present our method’s overarching simple learning algorithm.
@@ 3.1 Problem Setup
3.2 Algorithm
We describe the simple self-improving training scheme in Algorithm 1. We generate a set of random problem instances in each epoch and use the current best policy network to sample multiple trajectories for each instance. We add each instance and the corresponding best trajectory to the training dataset, from which the network is trained by imitation of the trajectories using a cross-entropy loss. At the end of the epoch, if the policy performs better when rolled out greedily than the currently best, we empty the dataset again (as we expect the sampled solutions to improve). Otherwise, the dataset will expand in the next epoch, providing the model with more training data.
@@ 3.2 Algorithm
Randomly initialize policy $\pi_{\theta}$ $\pi_{\text{best}}\leftarrow\pi_{\theta}$  normal-$\triangleright$ current best policy
1 $\textsc{Dataset}\leftarrow\emptyset$ foreach epoch do
2       Sample set of $n$ problem instances $\textsc{Instances}\sim\mathcal{X}$ foreach $x\in$ Instances do
             Sample set of $m$ feasible solutions $A:=\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(m)}\}\sim\pi_{\text{best}}$ $\textsc{Dataset}\leftarrow\textsc{Dataset}\cup\{\left(x,\operatorname*{arg\,max}_{\bm{a}_{1:T}\in A}f_{x}(\bm{a}_{1:T})\right)\}$   normal-$\triangleright$ add best solution
3
4      $\pi_{\theta}\leftarrow$ train supervised for one epoch with Dataset if greedy performance of $\pi_{\theta}$ on Validation better than $\pi_{\text{best}}$ then
             $\pi_{\text{best}}\leftarrow\pi_{\theta}$  normal-$\triangleright$ update best policy
             $\textsc{Dataset}\leftarrow\emptyset$  normal-$\triangleright$ reset dataset
5
6
@@ 2       Sample set of $n$ problem instances $\textsc{Instances}\sim\mathcal{X}$ foreach $x\in$ Instances do
The efficiency of the training is determined by the sampling method in line 7, highlighted in bold. The goal is to obtain a good solution with few samples $m$. Naively sampling sequences with replacement (i.e., Monte Carlo i.i.d. sampling, choosing the next token $a_{t}$ with probability $\pi(a_{t}|\bm{a}_{1:t-1})$) can lead to a homogeneous (often differing in a single token only (Li et al., 2016; Vijayakumar et al., 2018)) solution set with many duplicates. For example, Shi et al. (2020) show on a pre-trained TSP attention model (Kool et al., 2019b) that sampling 1280 sequences with replacement leads to $\sim$88duplicate sequences for instances with 50 nodes, and still $\sim$17duplicate sequences for instances with 100 nodes.
4 Sample without replacement, but with improvement
In this section, we describe our main contribution, a way to sample sequences without replacement in multiple rounds as described in Shi et al. (2020), coupled with a policy update after each round. We begin with a recall of Stochastic Beam Search (Kool et al., 2019c) using the Gumbel Top-k trick (Gumbel, 1954; Yellott Jr., 1977; Vieira, 2014), followed by a description of how to split the process into multiple rounds using an augmented trie, as described by Shi et al. (2020). We then describe how to combine the trie with a policy update. In the following, we use the term "beam search" to refer to the variant of beam search common in natural language processing (NLP), where nodes in the trie correspond to partial sequences, and a beam search of width $k$ keeps the top $k$ most probable sequences at each step, ranked by the total log-probability $\log\pi(\bm{a}_{1:t})=\sum_{t^{\prime}=1}^{t}\log\pi(a_{t^{\prime}}|\bm{a}_{1:t^{\prime}-1})$.
@@ 4.1 Preliminaries
Shi et al. (2020) present an incremental view on sampling sequences without replacement that can draw one sequence after another. They suggest maintaining an augmented trie structure that grows with the sampling process, where nodes correspond to partial sequences as in beam search. Starting from the root node (representing an empty sequence), a single leaf node $\bm{a}_{1:T}$ (representing a complete sequence) is sampled from the sequence model $\pi$. Then the total probability of $\pi(\bm{a}_{1:T})$ is recursively subtracted from the probability of the leaf and its ancestors (which correspond to all partial sequences $\bm{a}_{1:0},\bm{a}_{1:1},\dots,\bm{a}_{1:T}$). After renormalization of the probabilities, the trie represents a sequence model from which $\bm{a}_{1:T}$ has been removed, and re-sampling from this updated trie is equivalent to sampling from the model conditioned on the fact that $\bm{a}_{1:T}$ cannot be selected (see Figure 1). The incremental nature of the process has the advantage that we can sample continuously without replacement, stopping only when some condition is met. To further exploit the parallelism advantages of SBS, Shi et al. (2020) suggest combining SBS with their method: Instead of drawing just one sequence, we use the same trie for SBS and draw a batch of $k$ samples in parallel. Then, the recursive probability update is done for all $k$ sequences. This allows us to sample batches in rounds, e.g., with a beam width of $k$ and $n$ rounds, we sample $k\cdot n$ sequences without replacement. Note that although we must maintain a trie structure, each round of SBS maintains the constant memory requirements of a beam search of width $k$.
@@ 4.2.1 Trie update
We illustrate an example of the trie update after each round in Figure 2. We note that as all expectations are obtained via (4), the update (6) adds no significant computational overhead to the round-based SBS. When samling another $k$ sequences in the next round, we repeat the process.
@@
The update (6) scales the probabilities by $\exp(\sigma\sum A_{\bm{b}})$, so the role of the hyperparameter $\sigma$ is mainly to dampen the advantage. Since the magnitude of $\sum A_{\bm{b}}$ usually also changes as the model $\pi$ evolves, the choice of $\sigma$ is highly problem-dependent. In particular, one could think of, e.g., transforming the $A_{\bm{b}}$ by min-max normalization or changing $\sigma$ during training to optimize the behavior of the update. However, our experiments use the pure advantages and a constant, small problem-dependent $\sigma$ throughout the training process (see Section 5.3).

### v2

@@ 1 Introduction
Combinatorial Optimization (CO) plays a critical role in many real-world applications in fields as diverse as logistics, manufacturing, genomics and synthetic biology (Sbihi & Eglese, 2010; Crama, 1997; Naseri & Koffas, 2020). In a CO problem, the goal is to find the best solution from a finite set of options that maximizes a given objective function. The NP-hard nature of these problems and their complex variations make them exceptionally difficult to solve. Traditional methods for solving CO problems typically rely on exact algorithms, heuristics, and metaheuristics based on decades of research. Exact algorithms such as enumeration, cutting plane, and branch-and-bound solve CO problems to optimality, but are limited by NP-hardness (Laporte, 1992). On the other hand, (meta-)heuristic algorithms such as local search, genetic algorithms, and ant colony optimization can provide solutions faster without guaranteed quality (Crama et al., 2005; Yang, 2010; Dorigo et al., 2006). While powerful, the traditional methods struggle with scalability, computational efficiency, and adaptability to other problems. In addition, designing algorithms for a CO problem requires expert intuition and significant domain knowledge (Vesselinova et al., 2020).
To overcome these limitations, the success of deep learning has led to the emergence of Neural Combinatorial Optimization (NCO), which departs from traditional CO methods to leverage the pattern recognition and generalization capabilities of neural networks (Bengio et al., 2021). NCO aims to build models that can generate approximate solutions to CO problems by learning from data without manually crafting algorithmic rules.
@@
A prominent paradigm in NCO is the constructive approach, where a solution to a problem is built step by step in a sequential decision-making process. A neural network models the policy that guides these incremental decisions. This policy network is typically trained either by supervised learning (SL) (Vinyals et al., 2015; Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022; Hottung et al., 2020; Drakulic et al., 2023; Luo et al., 2023) or by reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019b; Nazari et al., 2018; Chen & Tian, 2019; Ahn et al., 2020; d O Costa et al., 2020; Hottung & Tierney, 2020; Kwon et al., 2020; Ma et al., 2021; Kim et al., 2021; Park et al., 2021; Wu et al., 2021; Park et al., 2022; Zhang et al., 2020; Hottung et al., 2022; Zhang et al., 2024). SL-based methods use expert solutions as labeled data for behavior cloning. While this training scheme is simple and effective, obtaining sufficient high-quality solutions from (exact) solvers can be expensive or even impossible for complex problems. On the other hand, RL methods exploit the natural modeling of the sequential problem as a finite Markov decision process, where the objective function is used as an episodic reward to maximize. In contrast to SL methods, they do not require pre-generated solutions but can suffer from strong hyperparameter sensitivity (Schulman et al., 2017; Henderson et al., 2018) and sparse rewards. State-of-the-art RL methods for NCO perform remarkably on the training distribution but generalize poorly to larger instances (Kwon et al., 2020; Luo et al., 2023): typically, the policy is trained with REINFORCE (Williams, 1992), where gradients are accumulated over full trajectories. This results in high memory requirements; hence, the autoregressive decoder of the underlying architecture is usually lightweight. Luo et al. (2023) and Drakulic et al. (2023) recently attributed the poor generalization ability to exactly the lightweight decoder of the used architectures (Drakulic et al., 2023; Luo et al., 2023). They show that increasing the size of the decoder structure and moving towards a decoder-only architecture addresses this issue and leads to superior generalization performance. However, the network’s size makes it impractical to train the policy with the commonly used RL methods.
To be able to train these larger architectures without expert data, we propose in this paper a simple, problem-independent training scheme on the intersection of RL and SL: We decode a set of solutions for randomly generated problem instances in each epoch using the current model. The solution with the best objective function evaluation for each instance is treated as a pseudo-expert solution (corresponding to a sequence of incremental decisions). Like decoder-only models in natural language processing (NLP), we train the policy on these pseudo-expert solutions in a next-token prediction manner. By repeating this process, we create a ’self-improving’ loop. While self-improving models by training them on their own output have been proposed in the context of NLP (Huang et al., 2023), our training strategy constitutes a novel proposition for NCO.
@@
The effectiveness of this strategy relies on the quality of the decoded solutions and the efficiency of the underlying sampling process. In particular, we want to sample as few sequences as possible to get better and better solutions in each epoch. While repeated sampling from the conditional distribution given by the model at each sequential step works in principle, it can result in a set of sequences that lacks diversity, contains numerous duplicates, and requires a large number of samples to find a solution that enables the network to improve. In this paper, we propose a sampling mechanism based on batch-wise Stochastic Beam Search (SBS) (Kool et al., 2019c) as introduced by (Shi et al., 2020). The mechanism draws samples without replacement in multiple rounds while maintaining a search tree. We suggest taking advantage of the batch- and round-wise mechanism to enhance its effectiveness. Given the sequences sampled in a single round, we estimate the expected value of the objective function and update the trie with the advantage of the sampled sequences, a strategy derived from a provable policy improvement operation. Furthermore, to balance the explore-exploit tradeoff, we couple this update strategy with Top-$p$ (nucleus) sampling (Holtzman et al., 2020). After each round, we gradually increase $p\leq 1$ to allow more unreliable sequences. Our method applies to any constructive neural CO problem, unlike sampling strategies that leverage problem specifics to diversify and improve the sampled solutions (Kwon et al., 2020).
@@
(i) We propose a novel ’self-improvement’ training strategy for NCO on the intersection of RL and SL: In each epoch, we decode a set of solutions from the current policy and treat the best sequences found as pseudo-expert trajectories on which we train the policy in a next-token prediction manner.
(ii) For sequence decoding, we propose a novel method based on drawing the sequences in multiple rounds using batch-wise SBS. We develop a technique to make the underlying search tree more informed in each round by updating the sequence probabilities according to their advantage. Informally, we increase (decrease) the probability of better (worse) trajectories than expected. This update adds next to no computational overhead to the sampling method, and we derive this technique from a provable policy improvement.
(iii) We show on the TSP and CVRP that our self-improvement method achieves comparable results to direct training on (near) optimal expert trajectories.
(iv) We present a novel transformer-based architecture for the JSSP that outperforms the current state of the art when trained with our method.
@@ 2 Related Work
Self-improvement in neural CO
The cross-entropy method (De Boer et al., 2005; Rubinstein, 1999) is a derivative-free optimization technique for a single problem instance that can be summarized as generating a set of solution candidates, evaluating them according to some objective function, and selecting the best-performing candidates to guide the generation of new solutions. While this principle of pushing the policy toward better performing sampled solutions lies at the heart of the self-critical policy gradient methods, there is little work on directly cloning the behavior of trajectories (from multiple instances) sampled from the current model in neural CO (Bogyrbayeva et al., 2022; Bengio et al., 2021). In fact, independent of our work, Corsini et al. (2024) make the same observation. They present a ’self-labeling’ strategy that samples multiple solutions from the current policy and uses the best one as a pseudo-label for supervised training. However, they focus only on the JSSP and use vanilla sampling with replacement, in contrast to our main contribution, which updates the policy during the sampling without replacement process. We refer to Appendix E for a deeper comparison. Furthermore, Luo et al. (2023) show a hybrid problem-specific way to learn without any labeled data by pre-training a randomly initialized model with RL on small instances, followed by reconstructing partial solutions to refine an unlabeled training dataset. We show that we can outperform this strategy without relying on RL.
@@ 3 Learning Algorithm
In the following section, we formally set up the problem and present our proposed overarching self-improvement learning method.
@@ 3.1 Problem Setup
3.2 Self-Improvement
We describe our proposed simple self-improving training scheme in Algorithm 1. We generate a set of random problem instances in each epoch and use the current best policy network to sample multiple trajectories for each instance. We add each instance and the corresponding best trajectory to the training dataset, from which the network is trained by imitation of the trajectories. In particular, we train the policy to predict the next token from a partial solution using a cross-entropy loss. At the end of the epoch, if the policy performs better when rolled out greedily than the currently best, we empty the dataset again (as we expect the sampled solutions to improve). Otherwise, the dataset will expand in the next epoch, providing the model with more training data.
@@ 3.2 Self-Improvement
1 Randomly initialize policy $\pi_{\theta}$
$\pi_{\text{best}}\leftarrow\pi_{\theta}$
  $\triangleright$ current best policy
2 $\textsc{Dataset}\leftarrow\emptyset$
3 foreach epoch do
4       Sample set of $n$ problem instances $\textsc{Instances}\sim\mathcal{X}$
5       foreach $x\in$ Instances do
6             Sample set of $m$ feasible solutions $A:=\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(m)}\}\sim\pi_{\text{best}}$
             $\textsc{Dataset}\leftarrow\textsc{Dataset}\cup\{\left(x,\operatorname*{arg\,max}_{\bm{a}_{1:T}\in A}f_{x}(\bm{a}_{1:T})\right)\}$
              $\triangleright$ add best solution
7
8      foreach batch do
9             Sample $B$ instances and partial solutions $\bm{a}_{1:d_{j}}^{(j)}=(a_{1}^{(j)},\dots,a_{d_{j}}^{(j)})$ with $d_{j}<T$ from Dataset for $j\in\{1,\dots,B\}$
             Minimize batch-wise loss $\mathcal{L}_{\theta}=-\frac{1}{B}\sum_{j=1}^{B}\log\pi_{\theta}\left(a^{(j)}_{d_{j+1}}|\bm{a}^{(j)}_{1:d_{j}}\right)$
              $\triangleright$ next-token prediction
10
11      if greedy performance of $\pi_{\theta}$ on Validation better than $\pi_{\text{best}}$ then
             $\pi_{\text{best}}\leftarrow\pi_{\theta}$
              $\triangleright$ update best policy
             $\textsc{Dataset}\leftarrow\emptyset$
              $\triangleright$ reset dataset
12
13
@@ 4       Sample set of $n$ problem instances $\textsc{Instances}\sim\mathcal{X}$
The efficiency of the training is determined by the sampling method in line 7, highlighted in bold. The goal is to obtain a good solution with few samples $m$, balancing exploration and exploitation. Naively sampling sequences with replacement (i.e., Monte Carlo i.i.d. sampling, choosing the next token $a_{t}$ with probability $\pi(a_{t}|\bm{a}_{1:t-1})$) can lead to a homogeneous (often differing in a single token only (Li et al., 2016; Vijayakumar et al., 2018)) solution set with many duplicates. For example, Shi et al. (2020) show on a pre-trained TSP attention model (Kool et al., 2019b) that sampling 1280 sequences with replacement leads to $\sim$88duplicate sequences for instances with 50 nodes, and still $\sim$17duplicate sequences for instances with 100 nodes.
4 Sample Without Replacement, but Improvement
In this section, we describe our main contribution, a way to sample sequences without replacement in multiple rounds as described in Shi et al. (2020), coupled with a policy update after each round. We begin with a recall of Stochastic Beam Search (Kool et al., 2019c) using the Gumbel Top-k trick (Gumbel, 1954; Yellott Jr., 1977; Vieira, 2014), followed by a description of how to split the process into multiple rounds using an augmented trie, as described by Shi et al. (2020). We then describe how to combine the trie with a policy update. In the following, we use the term "beam search" to refer to the variant of beam search common in NLP, where nodes in the trie correspond to partial sequences, and a beam search of width $k$ keeps the top $k$ most probable sequences at each step, ranked by the total log-probability $\log\pi(\bm{a}_{1:t})=\sum_{t^{\prime}=1}^{t}\log\pi(a_{t^{\prime}}|\bm{a}_{1:t^{\prime}-1})$.
@@ 4.1 Preliminaries
Shi et al. (2020) present an incremental view on sampling sequences without replacement that can draw one sequence after another. They suggest maintaining an augmented trie structure that grows with the sampling process, where nodes correspond to partial sequences as in beam search. Starting from the root node (representing an empty sequence), a single leaf node $\bm{a}_{1:T}$ (representing a complete sequence) is sampled from the sequence model $\pi$. Then the total probability of $\pi(\bm{a}_{1:T})$ is recursively subtracted from the probability of the leaf and its ancestors (which correspond to all partial sequences $\bm{a}_{1:0},\bm{a}_{1:1},\dots,\bm{a}_{1:T}$). After renormalization of the probabilities, the trie represents a sequence model from which $\bm{a}_{1:T}$ has been removed, and re-sampling from this updated trie is equivalent to sampling from the model conditioned on the fact that $\bm{a}_{1:T}$ cannot be selected (see Figure 1).
The incremental nature of the process has the advantage that we can sample continuously without replacement, stopping only when some condition is met. To further exploit the parallelism advantages of SBS, Shi et al. (2020) suggest combining SBS with their method: Instead of drawing just one sequence, we use the same trie for SBS and draw a batch of $k$ samples in parallel. Then, the recursive probability update is done for all $k$ sequences. This allows us to sample batches in rounds, e.g., with a beam width of $k$ and $n$ rounds, we sample $k\cdot n$ sequences without replacement. Note that although we must maintain a trie structure, each round of SBS maintains the constant memory requirements of a beam search of width $k$.
@@ 4.2.1 Trie Update
We illustrate an example of the trie update after each round in Figure 2. We note that as all expectations are obtained via (4), the update (6) adds no significant computational overhead to the round-based SBS. When samling another $k$ sequences in the next round, we repeat the process. In Appendix C.1, we experimentally compare the theoretical update (1) with the practical update (6), and see that while both yield strong results, (6) is almost always more favorable in practice.
@@
The update (6) scales the probabilities by $\exp(\sigma\sum A_{\bm{b}})$, so the role of the hyperparameter $\sigma$ is mainly to dampen the advantage. Since the magnitude of $\sum A_{\bm{b}}$ usually also changes as the policy $\pi$ evolves, the choice of $\sigma$ is highly problem-dependent. In particular, one could think of, e.g., transforming the $A_{\bm{b}}$ by min-max normalization or changing $\sigma$ during training to optimize the behavior of the update. However, our experiments use the pure advantages and a constant, small problem-dependent $\sigma$ throughout the training process (see Section 5.3).